OpenStudy (abdumu16):

A radioactive material decays according to the function A(t) = 500e^(-.0245t), where A is the amount in grams at the time (t years). What is the half-life of this material? Round your answer to the nearest year.

OpenStudy (lgnd_88):

-0.1386 divided by 2 and rounded

OpenStudy (lgnd_88):

-0.1386/2=-0.0693 rounded to be 700 years. 700 years is the answer

OpenStudy (mhchen):

What he said ^^

OpenStudy (irishboy123):

\(\large A(t) = 500e^{-\color{blue}{0.0245t}} = 500 \left( \dfrac{1}{2} \right) ^{\dfrac{t}{\tau} } \) where \(\large \tau\) is the half-life \(\large e^{-0.0245 ~ t} = \left( \dfrac{1}{2} \right) ^{\frac{t}{\tau} } \) \(\large -0.0245 ~ t = \dfrac{t}{\tau} \ln \dfrac{1}{2} \) \(\large -0.0245 = -\dfrac{1}{\tau} \ln 2 \) \(\large \tau = \dfrac{\ln 2 }{0.0245} \approx 28.3 \) yrs